OpenStudy (anonymous):

Solve: 6x − 10 + 14 = 22

OpenStudy (ash2326):

We have \[6x-10+14=22\] \[6x+4=22\] Subtract both sides by 4 \[6x+4-4=22-4\]\[6x+\cancel 4-\cancel 4=22-4\] \[6x=18\] Divide both sides by 6, we get \[\frac{6x}{6}=\frac{18}{6}\] \[\frac{\cancel 6 x}{\cancel 6} =3\] we get \[x=3\]